OpenStudy (anonymous):

Find the volume of the region between the graph of f(xy)=36−x^2−y^2 and the xy plane. I'm not too sure how to set up the double integral-- but 36-x^2-y^2 is what we will integrate? right??

OpenStudy (anonymous):

the x will surely go from 0 till 6

OpenStudy (anonymous):

or maybe from -6 to 6

OpenStudy (anonymous):

ok I guess I have an idea

OpenStudy (anonymous):

the level surface of this is always a circle

OpenStudy (anonymous):

when we look at the xy plane it is a circle with center (0,0) and radius 6

OpenStudy (anonymous):

so x goes from -6 to 6 and y goes from

OpenStudy (anonymous):

would we then use x^2 + y^2 = r^2 to solve for y?

OpenStudy (anonymous):

\[-\sqrt{36-x ^{2}}\] to \[\sqrt{36-x ^{2}}\]

OpenStudy (anonymous):

hmm this cannot be good... damn it would mean that the integral is 0

OpenStudy (anonymous):

I'm going to think about it over dinner-- will be back in 20 min

OpenStudy (anonymous):

ok, I am going to bed soon